OpenStudy (anonymous):

The LCM of 5 and 6 is ____. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

30

OpenStudy (anonymous):

Simply list the multiples of each number (multiply by 2, 3, 4, etc.) then look for the smallest number that appears in each list.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Choose the number that both 5 and 6 can divide?

OpenStudy (dan815):

what is the smallest number that both 5 and 6 can divide

OpenStudy (anonymous):

I was thinking the same that you will add that only.. :P

OpenStudy (dan815):

all numbers are some product of primes 5=5 6=3*2 the smallest number therefore that 6 and 5 can divide is a number with the divisions 6 and 5 in it 3*2*5 <--- therefore these primes must exist

OpenStudy (anonymous):

1

OpenStudy (anonymous):

IS IT 1

OpenStudy (dan815):

Guys, lowest common multiple means, the multilpes of 5 and 6 whhich has the same multiple that is the lowest

OpenStudy (anonymous):

You are finding GCF not LCM..:)

OpenStudy (anonymous):

@dyamondsr4ever 1 is HCF or GCF, it is not LCM..

OpenStudy (dan815):

there is a way to look at it list the multilpes of 5 5,10,15,20,25,30,35.. multilpes of 6 6,12,18,24,30,36 what numbers are common the smallest common multiple??? is 30

OpenStudy (anonymous):

I CANT FIGURE ITOUT PLEASE HELP ME

OpenStudy (anonymous):

Find a number so that when you divide the number by 5 you get 6. and if you divide the number by 6 you get 5.

OpenStudy (anonymous):

30

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Two things are there.. A largest number that can divide both given numbers --------> HCF A smallest number that can be divided by both given numbers ---> LCM

OpenStudy (anonymous):

THX GUYS I UNDERSTAND!!!! THX @dan815 FOR GIVING AN EXAMPLE

OpenStudy (dan815):

welcome

OpenStudy (dan815):

there is a more efficient way of doing LCM which is to look at primes factors

OpenStudy (dan815):

5=5 6=3*2 there fore 3*2*5 is the lowest possible LCM for example if u look at 4 and 8 LCM here 4=2*2 and 8 =2*2*2 all it has is an extra 2 we write out the common prime factors 2*2 from 4 and another 2 from 8 so 2*2*2 is the smallest possible multiple in other words like we already know 8 is the smallest LCM of 4 and 8 since 4*2 = 8

OpenStudy (dan815):

one simple thing to note is if you have a number X and Y and the prime factors of X do not include Y then the smallest possible LCM must be X * Y

OpenStudy (dan815):

for example 50 =5*10=5*5*2, therefore suppose i say hey whats the LCM of 3 and 150 since 3 is not a prime factor of 50 then 50*3 =150 must be the LCM